OpenStudy (anonymous):

what is the differential equation of all circles along x-axis with radius 2

OpenStudy (unklerhaukus):

so first find the equation of all circles along x-axis with radius 2

OpenStudy (anonymous):

What will be the equation of the circle along x-axis?? Is it?? \[(x-a)^2 = r^2\]

OpenStudy (anonymous):

I think I am wrong... Not sure...

OpenStudy (anonymous):

Or it can be: \[(x-a)^2 + y^2 = r^2\]

OpenStudy (unklerhaukus):

that second one looks right @waterineyes

OpenStudy (unklerhaukus):

now r=2

OpenStudy (anonymous):

I also think so.. Because y coordinate of centre is 0 along x axis..

OpenStudy (maheshmeghwal9):

i also think this

OpenStudy (unklerhaukus):

so solve for y^2

OpenStudy (anonymous):

Use this formula also: \[(a-b)^2 = a^2 + b^2 - 2ab\]

OpenStudy (anonymous):

(h,0) r=2 \[(x-h)^{2} + y ^{2}= r ^{2}\]

OpenStudy (anonymous):

Now apply the square formula that I have written above..

OpenStudy (anonymous):

but can i integrate it now

OpenStudy (anonymous):

You are dealing with differential equations and not intergral equations.. You have to eliminate the constants like h...

OpenStudy (anonymous):

yeah by integration...

OpenStudy (anonymous):

i will remove the arbitrary constants

OpenStudy (anonymous):

im wrong i must differentiate it rather.

OpenStudy (anonymous):

with resprct to x

OpenStudy (anonymous):

Yes you must differentiate it... Yes by doing so remove the arbitrary constant.. Yeh with respect to x take the derivative...

OpenStudy (anonymous):

\[y \prime \left[(x-h)^{2}+y ^{2}=4 \right] \rightarrow 2(x-h)(1-0) + 2y y \prime = 0\]

OpenStudy (anonymous):

is this correct @waterineyes

OpenStudy (anonymous):

Very true.. But you know h is there now also.. So, you will take the derivative again to eliminate h... Just go for it..

OpenStudy (anonymous):

yeah right\[\left[ 2(x-h)+2y y \prime = 0 \right] \div 2 \rightarrow (x-h)+y y \prime = 0 \rightarrow 1- 0 + y y'' + y' y' = 0 \rightarrow 1 + y y'' + y'^{2} = 0 \]

OpenStudy (anonymous):

\[1 + y y'' + (y')^{2} = 0\]

OpenStudy (anonymous):

done is this right??

OpenStudy (anonymous):

Yes absolutely... So, this is the answer..

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Welcome dear..